# Gradient estimates of the Finslerian Allen-Cahn equation

### v1

@@ 5 Local gradient estimates on forward complete Finsler manifolds
$0\displaystyle 2w^{-\frac{2}{q}}\phi^{2}H+(2-6s)\phi^{2}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\geq-(6s-4)\phi H=$ 0 . (5.7)
Plugging (5.6) and (5.7) into (5.5) and setting $0<q\leq\frac{3}{2(1+\sqrt{\frac{N-\epsilon(N-n)}{\epsilon}})}$ give that
$\begin{split}B\phi H\geq&\frac{3(1-\epsilon)}{2q^{2}(N-\epsilon(N-n))}(\phi H)^{2}\\ &-\frac{2}{q}(\phi H)g_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w)-2AK(2R)\phi H.\end{split}$ (5.8)
@@
$-\frac{2}{q}\phi Hg_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w)\geq-\frac{2}{q}(\phi H)^{\frac{3}{2}}\frac{C_{1}}{R}.$ (5.9)
Hence combining (5.8) and (5.9) yields
$B+2AK(2R)+\frac{2C_{1}}{qR}(\phi H)^{\frac{1}{2}}\geq\frac{3(1-\epsilon)}{2q^{2}(N-\epsilon(N-n))}(\phi H).$ (5.10)
@@
$\begin{split}\phi H&\leq\frac{q^{2}[N-\epsilon(N-n)]}{1-\epsilon}\left(B+\frac{2(N-\epsilon(N-n))}{1-\epsilon}\frac{C_{1}^{2}}{R^{2}}+2AK(2R)\right)\\ &=\frac{q^{2}[N-\epsilon(N-n)]}{1-\epsilon}\left\{\frac{2C_{1}^{2}+C_{2}}{R^{2}}+2AK(2R)\right.\\ &\quad+\frac{2(N+1-\epsilon(N+1-n))}{1-\epsilon}\frac{C_{1}^{2}}{R^{2}}\\ &\quad\left.+\frac{C_{1}}{R}\left[\sqrt{K(2R)C(N,A)}\coth\left(R\sqrt{\frac{K(2R)}{C(N,A)}}\right)+C_{0}(K_{0},A)\right]\right\}.\end{split}$ (5.11)
@@
$\begin{split}&\quad\phi^{2}\frac{2Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\geq-2AK(2R)\phi^{2}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &\geq-\frac{2(1-\epsilon)(s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}\phi^{2}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}-\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)^{2}}\phi^{2}A^{2}K^{2}(2R),\end{split}$ (5.12)
@@
$(6s-2)\phi^{2}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\leq(3s-1)^{2}\phi^{2}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+C^{4}\phi^{2},$ (5.13)
@@
$\begin{split}-\frac{2}{q}\phi Hg_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w)\leq&\frac{4(1-\epsilon)(sq+s-1)}{(N-\epsilon(N-n))s^{2}q^{2}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}H\phi^{2}\\ &+\frac{(N-\epsilon(N-n))s^{2}}{4(1-s)(sq+s-1)}\phi H\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}\phi)}{\phi}.\end{split}$ (5.14)
@@
We plug (5.12)-(5.14) into (5.5) and notice the above inequality to get
$\begin{split}\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\phi H)^{2}}{s^{2}q^{2}}\leq&\left(\frac{(N-\epsilon(N-n))s^{2}C_{1}^{2}}{4(1-\epsilon)(sq+s-1)R^{2}}+B\right)(\phi H)\\ &+\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)^{2}}A^{2}K^{2}(2R)+C^{4},\end{split}$ (5.15)
where we have abandoned several nonnegative terms. Recall a result that if $a_{0}\leq a_{1}+a_{0}a_{2}$ for some $a_{0},a_{1},a_{2}\geq 0$, then $a_{0}\leq a_{2}+\sqrt{a_{1}}$. (5.15) implies that
$\begin{split}\phi H\leq&\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)}\cdot\\ &\left\{\frac{(N-\epsilon(N-n))s^{2}C_{1}^{2}}{4(1-\epsilon)(sq+s-1)R^{2}}+\frac{(2C_{1}^{2})}{R^{2}}+\frac{C_{2}}{R^{2}}\right.\\ &\left.\quad+\frac{C_{1}}{R}\left[\sqrt{K(2R)C(N,A)}\coth\left(R\sqrt{\frac{K(2R)}{C(N,A)}}\right)+C_{0}(K_{0},A)\right]\right\}\\ &+\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)}AK(2R)+sq\sqrt{\frac{N-\epsilon(N-n)}{2(1-\epsilon)}}C^{2}.\end{split}$ (5.16)
@@
Remark 5.2.
Theorem 1.2 follows by taking $s=2$ in Case 2. and $\epsilon=\frac{1}{2}$, $q=\frac{2}{27(N+n)}$ in Theorem 5.1.
@@
Proof of Theorem 1.3.
In the case of $u\leq 1$, if $M$ is a compact Finsler manifold, Theorem 1.1 provides that
$\frac{F^{2}(\nabla u)}{u^{2}}+\frac{2}{3}(1-u^{2})\leq 0,$ (5.17)
since $K=0$. If $M$ is a forward complete Finsler manifold without boundary, Theorem 1.2 also implies (5.17), by letting $R$ tend to the infinity. However, (5.17) with $u\leq 1$ means that $\nabla u=0$ and $01-u^{2}=$ 0 in pointwise sense. ∎
Acknowledge
The author is supported partially by the NNSFC (No. 12001099, 12271093).

### v2

@@ 5 Local gradient estimates on forward complete Finsler manifolds
$0\displaystyle 2w^{-\frac{2}{q}}\phi^{2}H+(2-6s)\phi^{2}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\geq-(6s-4)\phi H=$ 0 . (69)
Plugging (68) and (69) into (67) and setting $0<q\leq\frac{3}{2(1+\sqrt{\frac{N-\epsilon(N-n)}{\epsilon}})}$ give that
$\begin{split}B\phi H\geq&\frac{3(1-\epsilon)}{2q^{2}(N-\epsilon(N-n))}(\phi H)^{2}\\ &-\frac{2}{q}(\phi H)g_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w)-2AK(2R)\phi H.\end{split}$ (70)
@@
$-\frac{2}{q}\phi Hg_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w)\geq-\frac{2}{q}(\phi H)^{\frac{3}{2}}\frac{C_{1}}{R}.$ (71)
Hence combining (70) and (71) yields
$B+2AK(2R)+\frac{2C_{1}}{qR}(\phi H)^{\frac{1}{2}}\geq\frac{3(1-\epsilon)}{2q^{2}(N-\epsilon(N-n))}(\phi H).$ (72)
@@
$\begin{split}\phi H&\leq\frac{q^{2}[N-\epsilon(N-n)]}{1-\epsilon}\left(B+\frac{2(N-\epsilon(N-n))}{1-\epsilon}\frac{C_{1}^{2}}{R^{2}}+2AK(2R)\right)\\ &=\frac{q^{2}[N-\epsilon(N-n)]}{1-\epsilon}\left\{\frac{2C_{1}^{2}+C_{2}}{R^{2}}+2AK(2R)\right.\\ &\quad+\frac{2(N+1-\epsilon(N+1-n))}{1-\epsilon}\frac{C_{1}^{2}}{R^{2}}\\ &\quad\left.+\frac{C_{1}}{R}\left[\sqrt{K(2R)C(N,A)}\coth\left(R\sqrt{\frac{K(2R)}{C(N,A)}}\right)+C_{0}(K_{0},A)\right]\right\}.\end{split}$ (73)
@@
$\begin{split}&\quad\phi^{2}\frac{2Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\geq-2AK(2R)\phi^{2}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &\geq-\frac{2(1-\epsilon)(s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}\phi^{2}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}-\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)^{2}}\phi^{2}A^{2}K^{2}(2R),\end{split}$ (74)
@@
$(6s-2)\phi^{2}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\leq(3s-1)^{2}\phi^{2}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+C^{4}\phi^{2},$ (75)
@@
$\begin{split}-\frac{2}{q}\phi Hg_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w)\leq&\frac{4(1-\epsilon)(sq+s-1)}{(N-\epsilon(N-n))s^{2}q^{2}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}H\phi^{2}\\ &+\frac{(N-\epsilon(N-n))s^{2}}{4(1-s)(sq+s-1)}\phi H\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}\phi)}{\phi}.\end{split}$ (76)
@@
We plug (74)-(76) into (67) and notice the above inequality to get
$\begin{split}\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\phi H)^{2}}{s^{2}q^{2}}\leq&\left(\frac{(N-\epsilon(N-n))s^{2}C_{1}^{2}}{4(1-\epsilon)(sq+s-1)R^{2}}+B\right)(\phi H)\\ &+\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)^{2}}A^{2}K^{2}(2R)+C^{4},\end{split}$ (77)
where we have abandoned several nonnegative terms. Recall a result that if $a_{0}\leq a_{1}+a_{0}a_{2}$ for some $a_{0},a_{1},a_{2}\geq 0$, then $a_{0}\leq a_{2}+\sqrt{a_{1}}$. (77) implies that
$\begin{split}\phi H\leq&\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)}\cdot\left\{\frac{(N-\epsilon(N-n))s^{2}C_{1}^{2}}{4(1-\epsilon)(sq+s-1)R^{2}}+\frac{(2C_{1}^{2})}{R^{2}}+\frac{C_{2}}{R^{2}}\right.\\ &\left.\quad+\frac{C_{1}}{R}\left[\sqrt{K(2R)C(N,A)}\coth\left(R\sqrt{\frac{K(2R)}{C(N,A)}}\right)+C_{0}(K_{0},A)\right]\right\}\\ &+\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)}AK(2R)+sq\sqrt{\frac{N-\epsilon(N-n)}{2(1-\epsilon)}}C^{2}.\end{split}$ (78)
@@
Remark 4.
Theorem 2 follows by taking $s=2$ in Case 2. and $\epsilon=\frac{1}{2}$, $q=\frac{2}{27(N+n)}$ in Theorem 11.
@@
Proof of Theorem 3.
In the case of $u\leq 1$, if $M$ is a compact Finsler manifold, Theorem 1 provides that
$\frac{F^{2}(\nabla u)}{u^{2}}+\frac{2}{3}(1-u^{2})\leq 0,$ (79)
since $K=0$. If $M$ is a forward complete Finsler manifold without boundary, Theorem 2 also implies (79), by letting $R$ tend to the infinity. However, (79) with $u\leq 1$ means that $\nabla u=0$ and $01-u^{2}=$ 0 in pointwise sense. ∎
[…]
Acknowledgments
The author is very grateful to the reviewers for their careful review and valuable comments.
Declarations
  • National Nature and Science foundation of China (No.12001099, 12271093)
  • Authors have no conflict of interest.
  • Ethics approval is not applicable
  • Data sharing is not applicable